### v1

Abstract
We provide a methodology for decoupling the bulk gravitational field equations of braneworld black holes to suppress the bulk singularities. To achieve this, we apply a linear Minimal Geometric Deformation (MGD) with respect to a coupling constant $\alpha$ to the $4D$ Minkowski spacetime embedded in an extra dimension. This results in a gravitational decoupling into a system $\mathcal{A}$ with equations of motion of order $\alpha^{0}$ and a system $\mathcal{B}$, related to the so-called Quasi-Einstein equations of order $\alpha$. This methodology allows for the construction of a regular geometry everywhere. We outline the necessary constraints for eliminating singularities and provide a recipe for solving the equations of motion. Both the warp factor, the scalar field, and the potential obtained are smooth and free from Dirac delta singularities. A control parameter is introduced such that, in the limit $b\to 0$, the Randall-Sundrum (RS) setup is recovered, resulting in a transition from a thick brane to a thin brane. The asymptotic behavior of the curvature invariant $\lim_{y\to\pm\infty}R_{5D}(r,y)$ is positive near the de Sitter core (for small $r$), asymptotically negative for finite $r>r_{*}$, and asymptotically flat at the $4D$ boundary as $r\to\infty$. Although this work aims to suppress bulk singularities, it is expected that our methodology may be useful for future investigations related to the embedding of gravitational objects within other braneworld contexts.
I introduction
The recent detection of gravitational waves by LIGO and VIRGO, coupled with the first image of a supermassive black hole captured by the Event Horizon Telescope (EHT) Abbott et al. (2016, 2017); Akiyama et al. (2022), has provided experimental data on gravity in the universe. These experiments have allowed us to test gravitational theories beyond General Relativity. In this context, braneworld theories have been tested using data from these experiments in recent years Banerjee et al. (2021, 2022); Visinelli et al. (2018); Mishra et al. (2022). The Randall-Sundrum (RS) model Randall and Sundrum (1999a, b) offers a geometric explanation for why gravity is weaker than the other fundamental forces of nature.
The RS model involves the embedding of four-dimensional flat branes within a five-dimensional anti-de Sitter (AdS) spacetime. On the other hand, since the gravitational collapse of a black hole is described by the Schwarzschild metric, it is natural to study this $4D$ geometry localized on a brane embedded in an extra dimension, that is, in a brane-world setup. This problem was initially investigated in reference Chamblin et al. (2000), specifically addressing the so-called black string. However, when embedding a Schwarzschild black hole in an extra dimension within the Randall-Sundrum framework, physical singularities arise both at the origin of the radial coordinate and at the infinity of the extra dimension. This can be tested by evaluating the Kretschmann scalar.
$K_{5D}\sim 40k^{4}+16M^{2}\frac{e^{4ky}}{r^{6}}$ (1)
It is also important to mention that the confinement of $3$-branes implies the presence of a pathology in the equations of motion at the brane locations, characterized by Dirac delta-type singularities. One way to address this pathology associated with the Dirac delta in the action principle of a $3$-brane was proposed in reference Kehagias and Tamvakis (2001). This approach consists of replacing the Dirac delta with a scalar field $\phi$ that resides in this space and is coupled through the standard action. Notably, in this model, $\phi$ physically represents the material from which the brane is constructed. Thus, resulting in a thick brane configuration. The warp factor $A(y)$ is defined to ensure that its value, as well as its first and second derivatives evaluated at the center of the extra coordinate domain ($y=0$), varies smoothly and takes on finite values. This contrasts with the Dirac delta-type discontinuities present in the Randall-Sundrum model. Furthermore, in these models, a control parameter $b$ is introduced, which establishes a specific limit where the geometry and matter configuration recover the Randall-Sundrum framework. That is to say, a transition occurs from a thick brane to the thin brane of the RS model. See examples of recent applications in Hendi et al. (2020); Peyravi et al. (2023); Rosa et al. (2022).
It is worth noting that the approaches to studying the representation of black holes (BHs) in a braneworld setup are primarily based on two methodologies: one originating from the brane and the other from the bulk. In the former approach, four-dimensional black hole solutions are constructed by studying the equations of motion induced on the brane Shiromizu et al. (2000). Some attempts to suppress the central radial singularity using 4D RBHs can be found in reference Neves (2021). However, this methodology does not fully capture the influence of the bulk geometry on the physical and geometric behavior of the four-dimensional spacetime, and, among other limitations, it does not allow for a complete analysis of singularities in the extra dimension. On the other hand, the second approach constructs spacetimes directly from the bulk, with the brane equations derived from there. This work will focus on this latter perspective. See for example Chamblin et al. (2000); Kanti et al. (2002); Hirayama and Kang (2001); Kanti et al. (2003, 2003) and more recent applications in Nakas and Kanti (2021a, b); Nakas et al. (2024); Crispim et al. (2024a, b).
On the other hand, in recent years, the Gravitational Decoupling (GD) algorithm Ovalle (2017) has garnered significant attention for yielding new solutions of physical interest. To name just a few cases, this method has been used to generate rotating black holes Contreras et al. (2021), hairy black holes Ovalle et al. (2021), spherically symmetric black holes Ovalle et al. (2018), regular black holes Ovalle et al. (2023), stellar distributions Estrada and Tello-Ortiz (2018); Estrada and Prado (2019), Charged Anisotropic Spherical Solutions Sharif and Sadiq (2018), solutions in Lovelock gravity Estrada (2019), among others. This method was also studied for braneworld models in reference León and Sotomayor (2019), but using the aforementioned methodology, based on the induced equations of motion on the brane. Broadly speaking, this method can be explained as follows: Initially, we have a gravitational system $\mathcal{A}$ whose geometry and matter source are of order $\alpha^{0}$. This geometry is linearly deformed with respect to a dimensionless coupling constant $\alpha$. This geometric deformation is coupled to a matter source, also of order $\alpha$, giving rise to a gravitational system $\mathcal{B}$, whose equations of motion are of order $\alpha$ and are so-called as Quasi-Einstein equations. Consequently, the equations of motion are decoupled into equations of order $\alpha^{0}$ for $\mathcal{A}$ and equations of order $\alpha$ for $\mathcal{B}$. First, the Einstein equations are solved for $\mathcal{A}$, followed by a simpler set of motion equations for $\mathcal{B}$. Finally, the two solutions are combined to derive the complete solution for the entire system.
Since the method mentioned in the previous paragraph (GD) has been used to test various gravitational scenarios, and given the growing interest in recent years to test braneworld models with experimental data Banerjee et al. (2021, 2022); Visinelli et al. (2018); Mishra et al. (2022), it would be quite beneficial to provide a methodology analogous to GD from the bulk perspective. This would enable the exploration of different gravitational scenarios within a braneworld setup.
In this work, we present a methodology to decouple the equations of motion for braneworld models, using the approach based on the bulk equations of motion. The methodology introduced here is analogous but distinct from the one mentioned previously, corresponding to reference Ovalle (2017). In this work, the proposed methodology aims to eliminate the singularity at infinity in the extra coordinate generated by embedding a black hole within a braneworld framework. To achieve this, we will initially consider a gravitational system $\mathcal{A}$ associated with the $4D$ Minkowski seed spacetime embedded in an extra dimension, along with matter sources given by the scalar field and its potential. We will then linearly deform the 4D Minkowski geometry with respect to the coupling constant $\alpha$. This geometric deformation will be coupled to a bulk energy-momentum tensor of order $\alpha$, giving rise to a new gravitational system $\mathcal{B}$, whose equations of motion, as we will see below, are also of order $\alpha$. These equations are analogous to the so-called quasi-Einstein equations. Consequently, through the equations of order $\alpha$, we will suppress the singularity at the radial origin, while the equations of order $\alpha^{0}$ will handle the singularity at infinity in the extra dimension. We will outline the constraints that must be satisfied to eliminate the singularity present both at the origin of the radial coordinate and at infinity in the extra coordinate. Additionally, we will provide a recipe for solving our equations of motion. Furthermore, we will identify the limits at which a thick brane in the bulk transitions to a thin brane and analyze the physical characteristics of our model.
Although this work presents a methodology aimed at eliminating the singularity at infinity, it is expected that the provided methodology will serve as a basis for studying other future problems related to the embedding of gravitational objects in a braneworld setup. For more complex scenarios, this would likely require redefining the coupling constants and the matter sources coupled in the bulk.
I.1 The geometry and Invariants of curvature
We will consider the following line element:
$ds^{2}=e^{2A(y)}\hat{g}_{\mu\nu}dx^{\mu}dx^{\nu}+dy^{2}$ (2)
where
$\hat{g}_{\mu\nu}dx^{\mu}dx^{\nu}=-U(r)dt^{2}+\frac{dr^{2}}{U(r)}+r^{2}d\Omega^{2}$ (3)
where $x^{\mu}=(t,r,\theta,\phi)$.
For the line element (2), the corresponding expression for the Ricci scalar is:
$R_{5D}=-20({A}^{\prime})^{2}-8{A}^{\prime\prime}+e^{-2A(y)}\bar{R}_{4D},$ (4)
where the indicates derivation with respect to the extra coordinate ($z$), and where $\bar{R}_{4D}$ corresponds to the Ricci scalar of the four dimensional geometry whose line element is given by equation (3):
$\bar{R}_{4D}=-\frac{\partial^{2}U}{\partial r^{2}}-\frac{4}{r}\frac{\partial U}{\partial r}+\frac{2}{r^{2}}\left(1-U\right).$ (5)
On the other hand, the Kretschmann scalar is given by:
$K_{5D}=40({A^{\prime}})^{4}+16({A^{\prime\prime}})^{2}+32{A^{\prime\prime}}({A^{\prime}})^{2}-4e^{-2A(y)}({A^{\prime}})^{2}\bar{R}_{4D}+e^{-4A(y)}\bar{K}_{4D},$ (6)
where the four dimensional Kretschmann scalar $\bar{K}_{4D}$ corresponds to the four dimenional geometry, equation (3)
$\bar{K}_{4D}=\left(\frac{\partial^{2}U}{\partial r^{2}}\right)^{2}+\frac{4}{r^{2}}\left(\frac{\partial U}{\partial r}\right)^{2}+\frac{4}{r^{4}}(1-U)^{2}.$ (7)
Below, we will present a list of criteria that must be satisfied for our geometry to be considered regular. These criteria point out that both the curvature invariants and the equations of motion (which we will test below) must be regular everywhere.
II Our propose
II.1 Action and equations of motion
Our action is determined by
[…]
where $\alpha$ is a dimensionless constant and where $L_{B}$ is the bulk lagrangian Kanti and Tamvakis (2002); Binetruy et al. (2000). Further below, we will explore the role that the constant $\alpha$ plays in the equations of motion.
Applying the variational principle to this action (8), we obtain the following set of equations:
$R_{MN}-\dfrac{1}{2}g_{MN}R=\dfrac{1}{4M^{3}}\biggl{(}\partial_{M}\phi\partial_{N}\phi-g_{MN}\biggl{[}\dfrac{1}{2}(\partial\phi)^{2}+V(\phi)\biggl{]}-\alpha\cdot T_{MN}^{{(bulk)}}\biggl{)},$ (9)
$\dfrac{1}{\sqrt{-g}}\partial_{M}(\sqrt{-g}g^{MN}\partial_{N}\phi)=\dfrac{\partial V(\phi)}{\partial\phi},$ (10)
$\nabla_{M}T^{MN}_{(bulk)}=0$ (11)
where $T^{MN}_{(bulk)}$ corresponds to the stress–energy tensor for bulk matter. As we will discuss below, for our geometric distribution, the corresponding energy-momentum tensor in the bulk must be diagonal, which we will write as:
$(T^{M}_{N})_{(bulk)}(r,y)=\mbox{diag}(-\rho,p_{r},p_{t},p_{t},T^{y}_{y})$ (12)
The conservation equation (11) gives rise to:
$\frac{\partial}{\partial r}p_{r}+\frac{2}{r}\left(p_{r}-p_{t}\right)$ $=0$ (13)
$\frac{\partial T^{y}_{y}}{\partial y}+\frac{\partial A}{\partial y}\left(4T^{y}_{y}-2p_{r}-2p_{t}\right)$ $=0$ (14)
II.2 Gravitational Decoupling–like background for the bulk
To explain our methodology, in this section we will highlight some aspects of our Gravitational Decoupling-like background for the bulk.
  1. 1.
    Initially, we consider a gravitational system $\mathcal{A}$, whose geometry and matter source are of order $\alpha^{0}$. That is, for this system we consider the coupling constant $\alpha$ to be turned off, thus, $\alpha=0$.
    • The matter sources for $\mathcal{A}$ are associated with the second and third terms in action (8), which are related to the scalar field and the potential, respectively.
    • The geometry of $\mathcal{A}$ corresponds to a seed spacetime where $(\hat{g}_{\mu\nu})_{seed}=\eta_{\mu\nu}$ , i.e., $U(r)_{seed}=1$, representing a Minkowski space embedded in an extra dimension, as given by the line element in equation(2).
    As we will see below, the equations of motion at order $\alpha^{0}$, i.e., with the coupling constant $\alpha$ turned off, correspond to equations (16), (17) and (10).
  2. 2.
    We now proceed to turn on the coupling constant $\alpha$, giving rise to the gravitational system $\mathcal{B}$, as follows:
    • The seed geometry is linearly deformed with respect to a dimensionless coupling constant $\alpha$ as follows
      $U(r)_{seed}=1\to U(r)=1-\alpha\cdot\frac{m(r)}{r}\Rightarrow({g}_{\mu\nu})_{seed}=\eta_{\mu\nu}\to\hat{g}_{\mu\nu}$ (15)
      Thus, the function $m(r)/r$ represents a minimal geometric deformation (MGD) Ovalle (2017) of Minkowski spacetime, thus, $\lim_{\alpha\to 0}\hat{g}_{\mu\nu}=\eta_{\mu\nu}$. That is, turning on the constant $\alpha$, due to the minimal geometric deformation, $U(r)_{seed}$ becomes $U(r)$, and for our case, $({g}_{\mu\nu})_{seed}=\eta_{\mu\nu}$ transitions to $\hat{g}_{\mu\nu}$.
    • This geometric deformation is coupled to a bulk matter source, which is also of order $\alpha$, as can be seen in the last term of equation (9). This gives rise to a gravitational system $\mathcal{B}$, whose equations of motion, as we will see below, are of order $\alpha$. These equations are analogous to the so-called quasi-Einstein equations Ovalle (2017).
  3. 3.
    As we will see below, as a consequence of the previous steps, the system of equations of motion will be decoupled into two simpler systems: the system $\mathcal{A}$, whose equations of motion are of order $\alpha^{0}$, and the system $\mathcal{B}$, whose equations of motion, the so–called Quasi Einstein equations, are of order $\alpha$.
II.3 Equations of motion
As mentioned in the introduction, this work provides a way to decouple the equations of motion in order to eliminate the singularity present at infinity in the radial coordinate. In this way, the constant $\alpha$ allows the decoupling of the equations into a part that depends on $(r,y)$ and a part that depends solely on the coordinate $(y)$. This form allows us to solve each component of the equations of motion separately. Below, we will describe why this way of decoupling will lead to a convenient way to find a solution corresponding to the coupled system.
We impose, for simplicity, that $\phi=\phi(y)$ and that $V(\phi)=V(\phi(y))$.
As we can see, the equations of motion split in the following manner:
II.3.1 (y) dependent equations of $\alpha^{0}$ order
Following the indications from point 1 of the previous section, we initially set the coupling constant $\alpha$ turned off, using as embedded seed spacetime $(\hat{g}_{\mu\nu})_{seed}=\eta_{\mu\nu}$, i.e., $U(r)_{seed}=1$, which represents a Minkowski space embedded in an extra dimension. The matter sources are provided by the scalar field and the potential. Thus, the $y$ dependent equations of motion of order $\alpha^{0}$ are:
$-\frac{1}{2}\left(\frac{\partial\phi}{\partial y}\right)^{2}+V(\phi)=-24M^{3}\left(\frac{\partial A}{\partial y}\right)^{2},$ (16)
$\frac{1}{2}\left(\frac{\partial\phi}{\partial y}\right)^{2}+V(\phi)=-12M^{3}\left(\frac{\partial^{2}A}{\partial y^{2}}\right)-24M^{3}\left(\frac{\partial A}{\partial y}\right)^{2}$ (17)
The conservation equation associated with the matter sources of order $\alpha^{0}$, i.e., the scalar field and the potential, is given by equation (10). Since two of these three equations are independent Hendi et al. (2020), due to the Bianchi identities, it is not necessary to develop the latter to obtain the solution.
II.3.2 Quasi Einstein (r,y) dependent equations of $\alpha$ order
By turning on the constant $\alpha$ and following the indications from point 2 of the previous section, we obtain the quasi-Einstein equations dependent on $(r,y)$ at order $\alpha$.
$p_{r}=\frac{8e^{-2A(y)}}{r^{2}}M^{3}\frac{\partial m}{\partial r}$ (18)
$p_{t}=\frac{4e^{-2A(y)}M^{3}\frac{\partial^{2}m}{\partial r^{2}}}{r}$ (19)
and
$T^{y}_{y}=\frac{8e^{-2A(y)}}{r^{2}}M^{3}\frac{\partial m}{\partial r}+\frac{4e^{-2A(y)}}{r}M^{3}\frac{\partial^{2}m}{\partial r^{2}}$ (20)
Comparing equations (18), (19), and (20), it is evident that
$T^{y}_{y}=p_{r}+p_{t}$ (21)
However, this must be consistent with equation (14).Thus, by introducing condition (21) into equation (14):
$T^{y}_{y}=e^{-2A(y)}\bar{T}^{y}_{y}(r)$ (22)
In this way, it is straightforward to verify that the energy-momentum tensor in the bulk has the following structure:
$(T^{M}_{N})_{(bulk)}=e^{-2A(y)}\mbox{diag}\left(-\bar{\rho}(r),\bar{p}_{r}(r),\bar{p}_{t}(r),\bar{p}_{t}(r),\bar{T}^{y}_{y}(r)\right)=e^{-2A(y)}\bar{T}^{M}_{N}(r)$ (23)
Thus, for the imposed five-dimensional geometry, the trace condition from reference Alencar et al. (2024) must be satisfied, namely:
$2T^{y}_{y}=T^{\mu}_{\mu}\Rightarrow 2\bar{T}^{y}_{y}=\bar{T}^{\mu}_{\mu}$ (24)
On the other hand, the radial component of the bulk conservation equation (13) reduces to:
$\frac{\partial}{\partial r}\bar{p}_{r}+\frac{2}{r}\left(\bar{p}_{r}-\bar{p}_{t}\right)=0$ (25)
It is worth stressing that the fact that the conservation equation of the system at order $\alpha^{0}$ (equation 10) is independent of the conservation equations of the Quasi-Einstein equations at order $\alpha$ (equations 24 and 25) implies that there is no exchange of energy-momentum between the fluid of the seed geometry (given by the scalar field and the potential) and the matter fields of the bulk. Therefore, there is only a purely gravitational interaction between both systems, $\mathcal{A}$ and $\mathcal{B}$. This is a common feature of the gravitational decoupling method.
Below, we will outline a list of criteria to ensure that the geometry is regular.
II.4 A list of criteria for ensuring the regularity of the geometry
Related to the curvature invariants:
  1. 1.
    In order to eliminate the singularity located at the radial origin, the embedded geometry must be regular everywhere. In this work, the embedded geometry corresponds to a Regular Black Hole (RBH). Thus, both $\bar{R}_{4D}$ and $\bar{K}_{4D}$ must be finite at every location.
    Although various types of RBH exist in the literature, for the sake of simplicity, this work will focus on the embedding of a 4D RBH with a de Sitter core. Regarding this point, it is important to note that finding a $4D$ geometry $\hat{g}_{\mu\nu}$ that satisfies this condition is not novel in the literature, as there is an extensive range of geometries representing $4D$ regular black holes (RBHs). Some examples in references Hayward (2006); Dymnikova (1992); Spallucci and Smailagic (2017). The main characteristics of RBHs with a de Sitter core that will be useful for this work are presented in Appendix A.
    However, as mentioned, part of the novelty of this work lies in our goal to establish a consistent framework for solving the bulk equations of motion and the bulk energy-momentum tensor through the gravitational decoupling of the equations of motion, in order to suppress the singularities present in the bulk.
  2. 2.
    In order to eliminate the singularity located at infinity in the extra dimension, the previous condition must be satisfied, see equations (4) and (6), and additionally, looking the same equations, both the function $A(y)$ and its first and second derivatives, $A^{\prime}(y)$ and $A^{\prime\prime}(y)$, must have finite values at every location included the infinity of the extra coordinate $y$.
Some remarks regarding the equations of motion:
  • The regularity conditions described in point 1, which are related to the conditions that $m$ and $\bar{\rho}$ must satisfy (outlined in Appendix A), along with the regularity conditions 2 for ${A}$, ensure that the equations of motion of order $\alpha$ are free from singularities everywhere.
  • The regularity conditions mentioned in 2 for $A^{\prime}$ and $A^{\prime\prime}$ ensure that the right-hand side of the equations of motion of order $\alpha^{0}$ is free of singularities everywhere. This condition also implies that the left-hand side of these equations must be free of singularities.
II.5 Proposed recipe for solving our decoupled equations
Our gravitational decoupling background can be interpreted as follows: we have a system $\mathcal{A}$ of order $\alpha^{0}$, which is associated with a seed spacetime and is coupled to the scalar field and potential matter sources. Conversely, the MGD of a seed spacetime gives rise to a system $\mathcal{B}$, governed by the so-called Quasi-Einstein equations, which are of order $\alpha$ and are coupled to the bulk energy-momentum tensor fields. Therefore, we propose the following procedure for solving the system:
  1. 1.
    Solve the $y$-dependent equations of motion of order $\alpha^{0}$:
    1. (a)
      Choose as a test of proof a function $A(y)$ that satisfies all the criteria outlined in point 2.
    2. (b)
      Compute the values of $\phi(y)$ and $V\left(\phi(y)\right)$.
  2. 2.
    Solve the $(r,y)$-dependent equations of motion of order $\alpha^{1}$:
    1. (a)
      Pick up an already existing solution $m(r)$ from the literature that satisfies condition 1, which is associated with the values of $\bar{\rho},\bar{p}_{r},$ and $\bar{p}_{t}$. Some examples in references Hayward (2006); Dymnikova (1992); Spallucci and Smailagic (2017).
    2. (b)
      Substitute the values of $\bar{\rho},\bar{p}_{r},\bar{p}_{t}$, and $A(y)$ into the energy-momentum tensor corresponding to the bulk 23.
II.6 An example of solving the equations of motion
II.6.1 Solve the $y$-dependent equations of motion of order $\alpha^{0}$
As test of proof we propose the function:
$A(b,y)=-k\sqrt{\frac{y^{2}+b^{2}}{b^{2}y^{2}+c^{2}}}$ (26)
It is direct to check that:
$\lim_{b\to 0}A(y,b)=-k(b/c)|y|\sim\mbox{\,RS warp factor,}$, (27)
[…]
$\displaystyle\lim_{y\to\pm\infty}{A}^{\prime}(y,b)=0$ (29)
$\displaystyle\lim_{y\to\pm\infty}{A^{\prime\prime}}(y,b)=0$ (30)
here indicates derivation respect to $y$. In order that $-A(b,y)$ be an increasing function, and thus, $-A(b,y\to\infty)>-A(b,0)$ we impose that
$c>b^{2}$ (31)
A typical behavior of the warp factor is shown in Figure 1. It is regular, exhibits $\mathcal{Z}_{2}$ symmetry, has a peak at the symmetry axis $y=0$, and decreases as one moves away from it. In the limit $b\to 0$, its behavior resembles that of a thin RS brane. It is worth nothing that, for thick branes, the variation around the peak is smooth.
Figure 1: Warp factor
We can also verify that:
[…]
$\displaystyle\lim_{y\to 0}{A}^{\prime}(y,b)=0$ (33)
[…]
From equation (33), we can see that the first derivative of $A$ is zero, indicating a smoother variation at $y=0$ compared to the RS model, which behaves like a sawtooth at this point. On the other hand, the second derivative, equation (34), has a finite negative value, unlike the RS model, where it behaves like a Dirac delta, blowing up to negative infinity at this point. From this point forward, for simplicity, we will set $k=1$.
Subtracting equations (17) and (16)
$(\phi^{\prime})^{2}=-12M^{3}A^{\prime\prime}=-12M^{3}\frac{b^{2}\left(-c^{4}-2b^{6}y^{2}+2b^{2}c^{2}y^{2}+3c^{2}y^{4}+b^{4}\left(c^{2}-3y^{4}\right)\right)}{\left(\frac{b^{2}+y^{2}}{c^{2}+b^{2}y^{2}}\right)^{3/2}\left(c^{2}+b^{2}y^{2}\right)^{4}}$ (35)
In relation to the previous paragraph, we can observe in Figure 2 that the second derivative of our warp factor has a finite value in the vicinity of $y=0$, which results in a smooth function rather than a Dirac delta distribution. We can also observe that, as $b\to 0$, where the warp factor approaches the RS behavior, the behavior of $A^{\prime\prime}$ resembles that of a Dirac delta function. Thus, as we sill see in equation (35) $(\phi^{\prime})^{2}\sim-A^{\prime\prime}\neq-\delta(y)$.
Figure 2: $A^{\prime\prime}$
Unfortunately, this equation does not have an analytical solution. However, both branches of the solution can be observed numerically. As shown in Figure 3, the behavior is analogous to that of a kink scalar field Rosa et al. (2022). The vertical distance between the two asymptotes is comparable to the distance between the two vacua $(-v\approx\phi_{min},v\approx\phi_{max})$ of a kink. Thus, The field $\phi$ interpolates between $(-v,v)$ as it varies along the extra dimension.
Figure 3: $\phi(y)$
Adding equations (17) and (16)
$V=-6M^{3}\left(4(A^{\prime})^{2}+A^{\prime\prime}\right)=-\frac{6\left(\frac{4(b^{4}-c^{2})^{2}y^{2}(c^{2}+b^{2}y^{2})}{b^{2}+y^{2}}+\frac{b^{2}\left(-c^{4}-2b^{6}y^{2}+2b^{2}c^{2}y^{2}+3c^{2}y^{4}+b^{4}(c^{2}-3y^{4})\right)}{\left(\frac{b^{2}+y^{2}}{c^{2}+b^{2}y^{2}}\right)^{3/2}}\right)}{(c^{2}+b^{2}y^{2})^{4}}$ (36)
In Figure 4, we observe the behavior of $V(y)$ and $\phi$ versus $V$, respectively. It is evident that both the scalar field and the potential, along with their respective derivatives, have finite values around $y=0$, indicating that the scalar field is smooth. This contrasts with the matter sources featuring a Dirac delta in the Randall-Sundrum model.
Figure 4: First: $V(y)$. Second $\phi$ vs $V$
Due to the $\mathcal{Z}_{2}$ symmetry of our warp factor, we can observe from the graphs that the potential also exhibits this symmetry. There are two degenerate minima of $V(\phi)$ that occur at $\phi=\pm v$, which represent the preferred values for the ground state, or so-called vacuum state, of $\phi$. The scalar field $\phi$ could be $-v$ in one particular region of spacetime and $+v$ in another. This latter scenario corresponds to a domain-wall configuration, where the domain wall is the relatively small region that separates the $-v$ vacuum from the $+v$ vacuum.
About the brane limit for our warp factor
It is straightforward to check that, in the limit as $b\to 0$, the second and third terms of the action (8) behave as follows:
$\bar{S}\in S=$ $-\lim_{b\to 0}\int d^{4}x\int dy\ \sqrt{-g}\biggl{(}\dfrac{1}{2}(\partial\phi)^{2}+V(\phi)\biggl{)}$
[…]
$\sim$ $24M^{3}\int\sqrt{-g}d^{5}x\Lambda_{eff}+12M^{3}\sigma\int d^{4}x\sqrt{-g^{(4D)}}$ (37)
where the first term correspond to the bulk term Kehagias and Tamvakis (2001) and the second term correspond to the brane term. This, we recovered the RS set–up in the limit $b\to 0$. Thus, $b$ can be interpreted as a control parameter, such that in the limit $b\to 0$, the RS structure is recovered.
II.6.2 About the solution of the $(r,y)$-dependent Quasi Einstein equations of order $\alpha^{1}$:
As previously mentioned, the novelty of this work does not lie in finding a 4D RBH geometry, but rather in providing a methodology to find a warp factor $A(y)$ that suppresses the singularity present at infinity in the extra coordinate. Thus, by following step 2a) of the recipe in subsection II.5, Any chosen geometry $m,\bar{\rho},\bar{p}_{r},\bar{p}_{t}$, when substituted into the energy-momentum tensor (equation 23), along with the warp factor $A(y)$ selected in the previous section, will correspond to a solution of the Quasi-Einstein equations (19), (20), (24) and (25)
II.6.3 Ricci Scalar and regularity of the geometry :
The value of the Ricci scalar is given by:
$R_{5D}(r,y)=\bar{R}_{4D}(r)\exp\left(2\sqrt{\frac{b^{2}+y^{2}}{c^{2}+b^{2}y^{2}}}\right)+\frac{4\left(-\frac{5(b^{4}-c^{2})^{2}y^{2}(c^{2}+b^{2}y^{2})}{b^{2}+y^{2}}-\frac{2b^{2}(-c^{4}-2b^{6}y^{2}+2b^{2}c^{2}y^{2}+3c^{2}y^{4}+b^{4}(c^{2}-3y^{4}))}{\left(\frac{b^{2}+y^{2}}{c^{2}+b^{2}y^{2}}\right)^{3/2}}\right)}{(c^{2}+b^{2}y^{2})^{4}}$ (38)
It is easy to check that the asymptotic value as $y\to\pm\infty$ is given by:
$\lim_{y\to\pm\infty}R_{5D}(r,y)=\bar{R}_{4D}(r)\exp(2/b)$ (39)
In this way, the sign of the Ricci curvature invariant in this asymptotic limit depends exclusively on the sign of $\bar{R}_{4D}$ for a given value of the radial coordinate.
Thus, following the earlier proposal, $R_{5D}$ is finite everywhere in the bulk, including the limits $r\to 0$ and $y\to\pm\infty$. Similarly, it is direct to check the same for the five-dimensional Kretschmann scalar given by Equation (6), allowing us to assert that our five-dimensional geometry is regular everywhere.
As mentioned in Appendix A, $\bar{R}_{4D}(r\approx 0)>0$ in the 4D de Sitter core. Then, $\bar{R}_{4D}(r>r_{*})<0$, and at the radial asymptote, we have $\lim_{r\to\infty}\bar{R}_{4D}\to 0$. Thus, the asymptotic value in the extra coordinate of $\lim_{y\to\pm\infty}R_{5D}(r,y)$ is
$\lim_{y\to\pm\infty}R_{5D}(r,y)\sim\begin{cases}\text{Positive close to the radial origin at the dS core}\\ \text{Negative for }r>r_{*}\\ 0\text{ for the radial asymptote}\end{cases}$
Figure 5: Ricci scalar $5D$ behavior using $R_{4D}$ oh Hayward model. First close to the radial origin at dS core. Second: for finite $r>r_{*}$ where $R_{4D}<0$ and for $r\to\infty$ where $R_{4D}\to 0$.
The generic behavior of the 5D Ricci scalar as a function of the extra coordinate $y$ is illustrated in Figure (5) for different fixed locations of the radial coordinate. First, we observe that close to the radial origin at the dS core, $R_{5D}$ is always positive and approaches a positive asymptote. Secondly, for finite values of $r>r_{*}$, where $R_{4D}<0$, the value of $R_{5D}$ is positive in the vicinity of the symmetry axis $y=0$, then changes to negative at a finite value of $y$, reaching a minimum, and subsequently increases to approach a negative asymptote. Finally, at the boundary of the radial coordinate, as $r\to\infty$, where $R_{4D}\to 0$, the value of $R_{5D}$ is positive near the symmetry axis $y=0$, then changes to negative at a finite value of $y$, reaching a minimum, and subsequently increases to approach an asymptote of zero.
Although the issue of stability under perturbations is beyond the scope of this work and could be explored in the future, it is worth mentioning that the behavior where $R_{5D}$ asymptotically approaches zero in the extra coordinate has been studied in Reference Germán et al. (2016) to address the stability problem associated with scalar perturbations of tachyonic thick branes. See other scenarios where $\lim_{y\to\pm\infty}R_{5D}\to 0$ in Koley and Kar (2005).
III Discussion and Summary
In this work, we have presented a methodology based on the gravitational decoupling of the bulk equations of motion to eliminate the singularity at infinity in the extra coordinate, which typically arises due to the embedding of a 4D black hole in a braneworld setup.
To establish our methodology, we use the Minkowski spacetime embedded in an extra dimension as the $4D$ seed geometry, which is linearly deformed with respect to a coupling constant $\alpha$, referred to as Minimal Geometric Deformation (MGD). Thus, the gravitational system decouples in the following manner: a system $\mathcal{A}$ of order $\alpha^{0}$, which is coupled to the scalar field and potential matter sources. Conversely, the MGD of Minkowski space leads to the system $\mathcal{B}$, related with the so-called Quasi-Einstein equations, which are of order $\alpha$ and are coupled to the bulk energy-momentum tensor fields. Additionally, we have determined the properties that the bulk energy-momentum tensor must satisfy. It is important to note that both gravitational sectors $\mathcal{A}$ and $\mathcal{B}$ do not exchange energy and momentum; rather, there is only gravitational interaction between them. We will outline the constraints that must be satisfied to eliminate the singularity present both at the origin of the radial coordinate and at infinity in the extra coordinate. Additionally, we will provide a recipe for solving our equations of motion. Consequently, the equations of order $\alpha$ eliminate the singularity at the radial origin, while the equations of order $\alpha^{0}$ eliminate the singularity at infinity in the extra dimension.
We obtain a regular geometry whose warp factor exhibits $\mathcal{Z}_{2}$ symmetry, has a peak at the symmetry axis $y=0$, and decreases as one moves away from it. In the limit $b\to 0$, its behavior resembles that of a thin RS brane. It is worth noting that, for thick branes, the variation around the peak is smooth. The first derivative of the warp factor is zero at the symmetry axis $y=0$, indicating a smoother variation at this point compared to the RS model, which behaves like a sawtooth function. On the other hand, the second derivative of the warp factor has a finite negative value, in contrast to the RS model, where it behaves like a Dirac delta function, blowing up to negative infinity at this point. Thus, the pathology of the equations of motion associated with the presence of the Dirac delta function is alleviated. We can also observe that as $b\to 0$, where the warp factor approaches RS behavior, the behavior of $A^{\prime\prime}$ resembles that of a Dirac delta function.
We have numerically plotted the behavior of the scalar field in Figure 3, which possesses two branches, resembling the behavior of a kink scalar field Rosa et al. (2022). The vertical distance between the two asymptotes is comparable to the distance between the two vacua $(-v\approx\phi_{\text{min}},v\approx\phi_{\text{max}})$ of a kink. Thus, the field $\phi$ interpolates between $(-v,v)$ as it varies along the extra dimension. Both the scalar field and the potential, along with their respective derivatives, have finite values around $y=0$, indicating that the scalar field is smooth. This contrasts with the matter sources featuring a Dirac delta function in the Randall-Sundrum model. Due to the $\mathcal{Z}_{2}$ symmetry, the potential has two degenerate minima, which occur at $\phi=\pm v$ and represent the preferred values for the ground state, or vacuum state, of $\phi$. The scalar field $\phi$ could be $-v$ in one particular region of spacetime and $+v$ in another. This latter scenario corresponds to a domain-wall configuration, where the domain wall is the relatively small region that separates the $-v$ vacuum from the $+v$ vacuum.
We have analyzed the behavior of the curvature $R_{5D}$ as a function of the extra coordinate $y$ for fixed values of the radial coordinate. We have determined that the asymptotic value of $R_{5D}$ at $y$ depends on the value of $\bar{R}_{4D}(r)$. Specifically, for values close to the de Sitter core located at the radial origin, the asymptote of $R_{5D}$ at $y$ is positive. For values of $r>r_{*}$, where $\bar{R}_{4D}(r)$ is negative, the asymptote of $R_{5D}$ at $y$ is negative. At the radial boundary of the geometry, where $\bar{R}_{4D}(r)$ transitions, the asymptote of $R_{5D}$ at $y$ is zero.
Although this work presents a methodology aimed at eliminating the singularity at infinity, it is expected that the provided methodology will serve as a basis for studying other future problems related to the embedding of gravitational objects in a braneworld setup. For more complex scenarios, this would likely require redefining the coupling constants and the matter sources coupled in the bulk.
As mentioned in the introduction, there has been significant interest in recent years in testing braneworld scenarios with experimental data Banerjee et al. (2021, 2022); Visinelli et al. (2018); Mishra et al. (2022). For example, reference Banerjee et al. (2022) demonstrated that, by embedding wormhole geometries in a braneworld scenario, the shadow data and the image diameter of Sgr A* fit better within braneworld models compared to the corresponding situation in General Relativity. Thus, from a theoretical point of view, it would be interesting if our methodology could serve as a basis for studying the embedding of other gravitational objects in a braneworld setup. For instance, one could investigate the gravitational decoupling related to the embedding of traversable wormholes, one-way wormholes, or regular black holes in a Simpson-Visser black bounce scenario Crispim et al. (2024a, b). This would require a suitable redefinition of the radial coordinate and an appropriate dependence of both this coordinate and the function $m$ on the coupling constants. This could be explored in future work
Acknowledgements
Milko Estrada is funded by ANID , FONDECYT de Iniciación en Investigación 2023, Folio 11230247. Tiago M. Crispim and Geová Alencar would like to thank Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and Fundação Cearense de Apoio ao Desenvolvimento Científico e Tecnológico (FUNCAP) for the financial support.
Appendix A Characteristics of the $4D$ RBH geometry
Typically, $4D$ RBH geometries exhibit the following characteristics:
  • The radial energy density, $\bar{\rho}$, reaches a maximum finite near the origin at $r=0$ and continuously decreases, satisfying $\lim_{r\to\infty}\bar{\rho}\to 0$. Consequently, the four-dimensional solution is asymptotically flat.
  • The mass function $m(r)$ behaves near the origin as $m\sim r^{3}$, and is a continuously increasing function. Since the $4D$ geometry is asymptotically flat, it approaches an asymptotic value, $\lim_{r\to\infty}m(r)=\bar{M}$, where $\bar{M}$ is a constant.
The two aforementioned points lead to:
  • The geometry behaves like a maximally symmetric space near the radial origin.
  • The geometry behaves like a flat Minkowski space at radial infinity.
It is worth noting that for our type of 4D RBH, the behavior of the Ricci scalar $\bar{R}_{4D}$ is generic, as illustrated in Figure 6, using the Dymnikova and Hayward models as examples. In this regard, we can highlight the following characteristics, which will be useful for analyzing the curvature of the entire 5D spacetime:
Figure 6: Ricci scalar $4D$
  • Near the radial origin, $r=0$, in the de Sitter core, $\bar{R}_{4D}$ has a positive maximum value.
  • From $r=0$, $\bar{R}_{4D}$ decreases, changing sign to negative at a point we will refer to as $r=r_{*}$, until it reaches a minimum value.
  • From this negative minimum, $\bar{R}_{4D}$ is a function that approaches $\bar{R}_{4D}\approx 0$ asymptotically.